# Concept-as-Tree: Synthetic Data is All You Need for VLM Personalization

### v2

@@ Abstract
Vision-Language Models (VLMs) have demonstrated exceptional performance in various multi-modal tasks. Recently, there has been an increasing interest in improving the personalization capabilities of VLMs. To better integrate user-provided concepts into VLMs, many methods use positive and negative samples to fine-tune these models. However, the scarcity of user-provided positive samples and the low quality of retrieved negative samples pose challenges for fine-tuning. To reveal the relationship between sample and model performance, we systematically investigate the impact of positive and negative samples (easy and hard) and their diversity on VLM personalization tasks. Based on the detailed analysis, we introduce Concept-as-Tree (CaT), which represents a concept as a tree structure, thereby enabling the data generation of positive and negative samples with varying difficulty and diversity for VLM personalization. With a well-designed data filtering strategy, our CaT framework can ensure the quality of generated data, constituting a powerful pipeline. We perform thorough experiments with various VLM personalization baselines to assess the effectiveness of the pipeline, alleviating the lack of positive samples and the low quality of negative samples. Our results demonstrate that CaT equipped with the proposed data filter significantly enhances the personalization capabilities of VLMs across the MyVLM, Yo’LLaVA, and MC-LLaVA datasets. To our knowledge, this work is the first controllable synthetic data pipeline for VLM personalization. The code is released at https://github.com/zengkaiya/CaT.
@@ 1 Introduction
[…]
Figure 1: The performance of Yo’LLaVA varies with the number of positive and negative samples. (Left) Various tasks show a noticeable decrease in performance when the number of positive samples is limited. (Right) As the number of retrieved negative samples increases, the performance of various tasks does not improve consistently. This might be due to uncertainty in the image retrieval process, leading to low-quality negative samples.
Figure 2: Overview of unified and controllable data synthesis pipeline and personalized model training. After a systematic analysis of positive and negative samples (Sec. 3.1), we utilize LLM and VLM to construct the concept tree and edit it to achieve controllable generation (Sec. 3.2). We then propose a well-designed data selection module and a new metric named PCS score to ensure the quality of synthetic data (Sec. 3.3). The ultimate high-quality data can be used to enhance test-time finetuning methods.
Although test-time fine-tuning has demonstrated its effectiveness, its success largely hinges on the availability of both positive and negative samples during the fine-tuning process. In real-world applications [53, 58, 8], positive samples available for personalizing models are often limited. For instance, users typically provide only 1 to 3 concept images rather than more than 10 for convenience, significantly limiting the model’s personalization performance. Furthermore, acquiring negative samples often depends on image retrieval [43], which is a cumbersome and unpredictable process, leading to difficulties in ensuring the quality of these negative samples. We performed preliminary experiments and the results in Fig. 1 support these findings. Although synthetic data can solve the issues mentioned, three major challenges still need to be addressed:
@@
    The specific roles of positive and negative samples in VLM personalization tasks are still not well understood. For instance, is the use of easy negative samples always necessary?
@@
Based on these findings, we propose Concept as Tree (CaT), a unified and controllable synthetic data framework for VLM personalization. Specifically, we define a three-layer tree as a representation of concept information in which the root node represents the concept category and the leaf nodes represent its attributes. By leveraging the power of Large Language Models and VLMs, we can automate the construction of trees for each concept. To synthesize positive samples, we use the root node as a prompt in the diffusion model, which is fine-tuned with concept images provided by the user. To create easy negative samples, we modify the root node information and generate a tree for controllable data generation. To create hard negative samples, we modify the leaf nodes of the concept-correlated tree in order to generate the prompt. Controlling diversity is equivalent to varying the number and types of editing operations performed on the tree during each generation.
A comprehensive data synthesis process must ensure the quality of generated data. Therefore, we propose a well-designed and easily implementable method for data filtering. We state that the information in an image includes two types: concept-specific features that are unique to each concept and concept-agnostic features that possess general characteristics [40]. Therefore, we apply the same perturbation to the synthesized images and calculate the distance between the synthesized images and the user-provided images both before and after the perturbation. The difference between these two distances can be defined as a Perturbation-based Concept-Specific (PCS) score, which serves as a key metric for assessing sample quality. We set filtering thresholds based on PCS score to ensure high data quality when filtering various types of generated samples. Through the aforementioned methods, we achieved a unified and controllable pipeline of data synthesis, as shown in Fig 2.
@@ 2 Related work
Personalization for Vision Language Models. Vision-Language Models(VLMs) [38, 64, 22] have exhibited remarkable capabilities across diverse domains, such as data mining [42, 28], fine-grained understanding [36, 25], and visual question answering [7, 10]. To seamlessly integrate these models into daily life, there has been growing interest in VLM personalization [2, 44, 4, 47]. This task is first introduced by MyVLM [2], employing an additional module strategy for improved injection of concept information into VLMs. YoLLaVA [44] and MC-LLaVA [4] utilized efficient end-to-end fine-tuning methods to tackle challenges in both single and multi-concept scenarios. The performance of these fine-tuning methods is highly dependent on the positive and negative samples. However, users often employ a limited number of concept images (i.e., 1 to 3 images), and the challenge of acquiring negative samples complicates fine-tuning approaches. In this study, we investigate the impact of positive and negative samples and their diversity on VLM personalization tasks. And we propose a complete data synthesis pipeline, thus overcoming challenges in training VLM models with limited data availability.
Learning from Synthetic Data. The use of synthetic data [31, 62, 20, 11] has been extensively studied across various computer vision tasks, such as classification [52], segmentation [59], and object detection [17]. With the rapid consumption of existing data resources driven by the development of Large Language Models (LLMs) and VLMs, researchers aim for models to achieve further improvements using synthetic data [45, 34, 41, 29]. Synthetic corpora are generated for instruction fine-tuning of large models. At the same time, large models leverage recaptioned datasets to enhance quality and improve CLIP performance [16]. To better support the learning of VLMs, direct generation of text-image pairs is pursued [35]. However, existing generation methods only consider class information [55]. These methods rely on simplistic prompt templates, attributes, and class information, either manually defined or LLM-generated [14, 56, 30]. To address the issue of insufficient data in VLM personalization scenarios, we supplement the dataset by synthesizing concept-centric positive and negative sample data. We adopt a unified framework to control the positivity and diversity of the generated samples.
[…]
@@ 3 Method
We present a comprehensive data synthesis pipeline to address the challenge of personalizing Vision-Language Models (VLM). This pipeline views the concept as a tree structure and generates both positive and negative samples through tree operations. It also encompasses a well-designed filtering strategy to select the generated samples. The overall design of our pipeline is illustrated in Fig. 2. Specifically, in Sec. 3.1, we systematically study the impact of positive and negative samples on VLM personalization tasks. Building on this, we introduce the CaT framework and use tree operations to generate synthetic positive and negative samples in Sec. 3.2. Furthermore, to ensure the quality of the synthetic samples, we propose a filtering strategy based on the PCS score in Sec. 3.3.
@@ 3.1 Impact of Positive and Negative Samples
Figure 3: Explore the role of positive and negative samples in personalization and their demand for diversity. (Left) To better investigate sample effectiveness, we conduct experiments on half the number of positive and negative sample contrasts to the original setting. Adding positive samples brings a general raising among the tasks while easy and hard negative samples have their different improvement preferences. (Middle&Right) Keeping the number of samples unchanged, we vary the diversity of easy and hard negative samples. The performance improvements brought by them vary across different tasks. However, hard negative samples are more sensitive to diversity. Moreover, excessive diversity may lead to a decline in task performance. The diversity score is calculated by clustering the retrieved negative samples using K-means and measuring the distance of each data point to the cluster centroids.
In this section, we examine the effects of positive and negative samples and their diversity on the personalization of VLM. Based on the positive samples provided by the MC-LLaVA dataset and negative samples retrieved online, we conducted a series of experiments with the Yo’LLaVA method. As illustrated in Fig. 3 left, we found that an increase in positive samples consistently improves outcomes across all tasks. While easy negative samples primarily enhance recognition abilities, hard negative samples improve conversational capabilities. Interestingly, the captioning task fundamentally reflects conversational abilities. However, since the evaluation method assesses the presence of a concept identifier related to recognition capabilities, the training data that includes both easy and hard negative samples yields the best performance. Furthermore, we find that easy negative samples are essential. Using only hard negative samples can help the model memorize the visual features of a concept and perform well in dialogue tasks. However, its ability to recognize the concept identifier significantly declines. Easy negative samples are irreplaceable, as they are diverse and widely sourced, providing valuable context for the model’s learning process. Recognition tasks are often the initial step for a VLM acting as a personal assistant, representing an essential capability that is crucial.
Based on the above mentioned observation, we conducted additional experiments that examined different levels of diversity for negative samples. Here, diversity is calculated by clustering a set of negative samples and calculating the distance from the cluster center to each sample. From Fig. 3 middle and right, keeping the number of positive samples unchanged, it is evident that the gain curves for easy and hard negative samples differ across various tasks, indicating that their diversity requirements are distinct. Low diversity of both easy and hard negative samples can simultaneously restrict the model performance trained on them. When the diversity of hard negative samples is excessively high, it often introduces noise during training, resulting in a decline in model performance. These experiments clarify the role of diversity and motivate us to manage the diversity when creating negative samples. Therefore, we propose representing the concept as tree to facilitate the generation of negative samples with various difficulties and diversities.
@@ 3.2 Concept as Tree Framework
Figure 4: Using concept trees to synthesize positive and negative samples and three editing operations for constructing diverse concept trees. (a) We first utilize the CaT framework and reference images to obtain the concept tree $T_{C}$. Then, we employ a fine-tuned diffusion model along with the information from the root node $R(C)$ to synthesize positive samples. For easy negative samples, we modify the class of the root node $R(C)$ and obtain a new concept tree for synthesis. Finally, we apply three editing operations to the original concept tree to obtain multiple trees for synthesizing hard negative samples with diversity. The model used for synthesizing negative samples does not require fine-tuning. (b) We visualize the three mentioned tree editing operations in detail. Compared with the reference image, adding a dimension of “mood” actually leads to different emotions in cats. After removing the “object” dimension, the environment around the cats becomes simpler. By modifying a dimension, the behavior of cats in the composite image becomes more diverse.
Concept Tree Representation and Construction. To enable controllable data generation, inspired by SSDLLM [42], we represent each concept $C$ as a hierarchical tree $T_{C}=(R(C),D_{C},A(D_{C}))$. As shown in Fig. 4(a), $R(C)$ serves as the root node, encapsulating the high-level category of the concept (e.g., “cat”, “dog”, etc.), while $D_{C}=\{D_{1},D_{2},\dots,D_{m}\}$ defines a set of attribute dimensions that distinguish various aspects (e.g., “appearance”, “behavior”, “location”, etc.). Each dimension $D_{i}$ is associated with a set of attributes $A(D_{i})$, collectively forming $A(D_{C})=\{A(D_{1}),A(D_{2}),\dots,A(D_{m})\}$ (for instance, the “behavior” dimension might include attributes like “sitting”, “lying” and “climbing”, etc.).
@@
Tree Editing Operations. We define three fundamental tree editing operations $F(R(C),D_{C}^{\prime})$ (see Fig. 4(b)), which modify $T_{C}$ to systematically generate negative samples:
@@ 3.3 Sample Filtering with PCS Score
Figure 5: Filtering method based on the PCS score and the visualization of the selection of high-quality images. (a) We perform patch shuffling on the synthetic images and extract the features $F_{o}$, $F_{d}$ and $F_{r}$ from the original images, disturbed images and reference images, respectively. Subsequently, we calculate the image similarities $S_{o}$ and $S_{d}$, then select high-quality images with high PCS score. (b) The reference images are marked with blue box, images containing more CS information are highlighted with green box and those containing more CA information are indicated with red box. Using image cosine similarity alone cannot distinguish between these two types of synthetic images, but our proposed PCS score can effectively filter out the images on the right.
To ensure the quality of synthesized samples, we implement a filtering process for all generated samples. We classify the information contained in images into two types: concept-specific (CS) and concept-agnostic (CA). The former refers to features that are distinctive to the concept, such as appearance or type, while the latter encompasses other elements present in the image, such as background or weather conditions. We consider synthetic data predominantly containing CS information to be regarded as high-quality samples. Conversely, samples with relatively more CA information may be categorized as low-quality samples.
Inspired by this, we first mix patches from the reference and original images, then calculate the similarity between the altered images and the reference images using the CLIP model [46]. The difference between the similarities before and after perturbation is defined as the Perturbation-based Concept-Specific (PCS) score. Typically, the PCS score of an image should be quite high because mixing patches alters the appearance or shape of objects within the image, thus disrupting the original CS information. This disruption causes CLIP to struggle to understand the conceptual relationship between the synthetic image and the reference image, leading to a notable reduction in the similarity of the altered image. However, if the similarity of an image relies not on CS information but on CA information, CLIP might perceive uncommon or incorrect object relationships with the reference image. In such cases, the similarity before and after perturbation changes little, meaning the PCS score will be small. Therefore, we claim that the PCS score can be used to determine the information components in the image. A high PCS score indicates that the image carries more CS information, while a low PCS score suggests that the image contains more CA information.
As illustrated in Fig. 5(a), let the reference image be $I_{\text{reference}}\in\mathcal{D}_{\text{user}}$, the synthetic image before perturbation be $I_{\text{original}}$ and the image after perturbation be $I_{\text{disturbed}}$. We extract the visual features $F_{r},F_{o}$ and $F_{d}$ of the corresponding images using CLIP [46] image encoder $f_{\theta}$:
@@
The difference between $S_{o}$ and $S_{d}$ is the PCS score. Based on this score, we can select high-quality synthetic data, i.e., $G(I)=\{I\mid\text{PCS}(I)>\tau_{PCS}\}$, where $G(I)$ represents high-quality samples and $\tau_{PCS}$ is the threshold for containing more CS information. As shown in Fig. 5(b), when only using the filtering method based on cosine similarity, both the images in the green and red boxes achieve high similarity scores. However, the concept in the red box lacks specific visual features of the reference image’s concept, such as hair, facial structure and body shape. Instead, the background and other CA information in these images have a greater impact on the similarity. However, through the newly proposed perturbation-based filtering method, we can effectively identify and filter out images with more CA information using PCS score, while retaining high-quality images with more CS information for better personalization.
4 Experiments
4.1 Experimental Setup
Implement Details. Regarding the experiments conducted on all datasets, the quantity of positive and negative samples for all the baselines follows their original settings. The training data used for Baseline (syn) is entirely generated by CaT. To ensure the consistency of the positive sample count, Baseline (Real+Syn) is articulated as comprising 1 to 3 original concept images and several synthesized positive samples. Baseline (Real+Syn) (Plus) indicates that an equal number of synthesized positive samples have been added to the previous positive samples. All negative sample quantities in the above experiments are strictly controlled for consistency. When the tree dimension corresponding to the concept is greater than or equal to 5, we do not use Add. When it is less than or equal to 3, we do not use Remove. For other instances, we utilize the three editing operations evenly for experiments. With regard to the synthesized images, we utilize GPT-4o to generate the instruction text pairs required for their training. Details of baseline, datasets and training hyperparameters can be found in the Appendix.
@@ 4.1 Experimental Setup
GPT-4o [1]+Prompt 0.746 0.888 0.712 0.728 0.836 0.856 0.932 0.897 0.891 0.969
MyVLM [2](Real) 0.795 0.779 - 0.640 0.714 0.911 0.897 - 0.938 0.921
@@
RAP-MLLM [23] 0.747 0.832 0.709 0.424 0.711 0.845 0.917 0.874 0.870 0.937
Table 1: Results of synthetic data used in different personalization methods. All methods achieve improvements across three datasets when using synthetic data. When trained with a combination of original positive data and all synthetic data, all methods can even achieve results comparable to GPT-4o. Noting, the Choice-T is a pure text task, but MyVLM requires concept images to load the corresponding embeddings during testing time, which makes it unable to perform the Choice-T task. The green numbers represent the improvement of the new SOTA compared to original baseline. Real = Original Data; Syn = Synthetic Data; P = Plus.

### v3

# Concept-as-Tree: A Controllable Synthetic Data Framework Makes Stronger Personalized VLMs

@@ Abstract
Vision-Language Models (VLMs) have demonstrated exceptional performance in various multi-modal tasks. Recently, there has been an increasing interest in improving the personalization capabilities of VLMs. To better integrate user-provided concepts into VLMs, many methods use positive and negative samples to fine-tune these models. However, the scarcity of user-provided positive samples and the low quality of retrieved negative samples pose challenges for existing techniques. To reveal the relationship between sample and model performance, we systematically investigate the amount and diversity impact of positive and negative samples (easy and hard) on VLM personalization tasks. Based on the detailed analysis, we introduce Concept-as-Tree (CaT), which represents a concept as a tree structure, thereby enabling the data generation of positive and negative samples with varying difficulty and diversity, and can be easily extended to multi-concept scenarios. With a well-designed data filtering strategy, our CaT framework can ensure the quality of generated data, constituting a powerful pipeline. We perform thorough experiments with various VLM personalization baselines to assess the effectiveness of the pipeline, alleviating the lack of positive samples and the low quality of negative samples. Our results demonstrate that CaT equipped with the proposed data filter significantly enhances the capabilities of VLMs across personalization benchmarks. To the best of our knowledge, this work is the first controllable synthetic data pipeline for VLM personalization. The code will be released.
@@ 1 Introduction
Vision-Language Models (VLMs) [bai2023qwenvl, lin2023sphinx, li2023blip] have produced impressive results across various tasks, showcasing their potential as AI assistants [dorka2024training, zhang2024vision, hong2025dialoguelanguagemodellargescale]. Despite their success, VLMs still struggle to generate personalized responses, such as including user-specific identifiers like $\langle\text{Bob}\rangle$ or $\langle\text{Lina}\rangle$ in the conversation. Several recent studies [alaluf2025myvlm, nguyenyo, an2024mcllavamulticonceptpersonalizedvisionlanguage, hao2024rememberretrievegenerateunderstanding] have explored the personalization of VLMs to address this challenge and seamlessly incorporate VLMs into everyone’s daily life. Among these, the test-time fine-tuning paradigm exemplified by Yo’LLaVA [nguyenyo] demonstrates simplicity and effectiveness, while preserving the models’ existing knowledge as much as possible.
Figure 1: The performance of Yo’LLaVA on Yo’LLaVA Dataset. (Left) Various tasks show a noticeable decrease in performance when the number of positive samples is limited. (Right) As the number of retrieved negative samples increases, the performance of various tasks does not improve consistently. This might be due to uncertainty in the image retrieval process, leading to low-quality negative samples. More results on different datasets can be found in Appendix.
Figure 2: Overview of unified and controllable data synthesis pipeline and personalized model training. After a systematic analysis of positive and negative samples, we utilize LLM and VLM to construct the concept tree and edit it to achieve controllable generation. We then propose a well-designed data selection module and a new metric named PCS score to ensure the quality of synthetic data. The ultimate high-quality data can be used to enhance test-time fine-tuning methods.
Although test-time fine-tuning has shown its effectiveness, its success largely hinges on the availability of both positive and negative samples during the process. In real-world applications [song2021bob, welch2022leveraging, chawla2013bringing], positive samples available for personalizing models are often limited. For instance, users typically provide only 1 to 3 concept images rather than more than 10 for convenience, significantly limiting the model’s personalization performance. Furthermore, the acquisition of negative samples often relies on image retrieval [meghwani2024enhancing, nguyenyo], a cumbersome and unpredictable process that complicates the assurance of their quality. Preliminary experiments, as illustrated in Fig. 1, corroborate the limitations. Given these challenges, a natural idea is to use synthetic data to provide controllable supplementation for both positive and negative samples. While this approach may help mitigate the aforementioned issues, three significant challenges remain to be addressed:
@@
    The roles for positive and negative samples in VLM personalization are still not well understood. For instance, is the use of easy negative samples always necessary?
@@
Based on these findings, we propose Concept as Tree (CaT), a unified and controllable synthetic data framework for VLM personalization. Specifically, we define a three-layer tree as a representation of concept information in which the root node represents the concept category and the leaf nodes represent its attributes. By leveraging the power of LLMs and VLMs, we can automate the construction of trees for each concept. To synthesize positive samples, we directly employ the root node as a prompt for diffusion model. Easy negative samples are generated by altering the root node information and constructing a tree for controllable data generation. Hard negative samples are produced according to concept tree whose leaf nodes are strategically adjusted. Moreover, as the number of concepts increases, we naturally merge the trees into forests according to the specific scenarios [an2024mc] (e.g., two concepts). Because of the designed tree structure, our framework supports all types of personalized data. Meanwhile, controlling diversity is equivalent to varying the number and types of editing operations performed on the tree in our framework.
Figure 3: Explore the role of positive and negative samples in personalization and their demand for diversity. (Left) We evaluate the effect of halving positive and negative samples. Positive samples generally improve performance, while easy and hard negatives show task-specific benefits. (Middle & Right) With fixed sample counts, we vary the diversity of easy and hard negatives. Hard negatives are more sensitive to diversity, and excessive diversity can degrade performance. Diversity scores are computed by clustering retrieved negatives via K-means and measuring distances to cluster centroids. The first line shows results on the Yo’LLaVA dataset; the second line illustrates results on the MC-LLaVa dataset.
A comprehensive data synthesis process must ensure the quality of generated data. Therefore, we propose a well-designed and easily implementable method for data filtering. We state that the information in an image includes two types: concept-specific features that are unique to each concept and concept-agnostic features that possess general characteristics [liu2023vida]. Therefore, we apply some perturbation to the synthesized images and calculate the distance between the synthesized images and the user-provided images both before and after the perturbation. The difference between these two distances can be defined as a Perturbation-based Concept-Specific (PCS) score, which serves as a key metric for assessing sample quality. We set filtering thresholds based on the PCS score to ensure high data quality when filtering various types of generated samples. Through the aforementioned methods, we achieved a unified and controllable pipeline of data synthesis, as shown in Fig. 2.
@@ 2 Related work
Personalization for Vision Language Models. Vision-Language Models(VLMs) [liu2023visual, li2024llava, li2024llava_next] have exhibited remarkable capabilities across diverse domains, such as data mining [luo2024llm], fine-grained understanding [lin2024draw], and visual question answering [cao2025move]. To seamlessly integrate these models into daily life, there has been growing interest in VLM personalization [alaluf2025myvlm, nguyenyo, an2024mc, hao2025rap, an2025unictokens]. This task is first introduced by MyVLM [alaluf2025myvlm], employing an additional module strategy for improved injection of concept information into VLMs. YoLLaVA [nguyenyo] and MC-LLaVA [an2024mc] utilized efficient end-to-end fine-tuning methods to tackle challenges in both single and multi-concept scenarios. The performance of these fine-tuning methods is highly dependent on the training samples. However, users often employ a limited number of concept images (i.e., 1 to 3), and the challenge of acquiring negative samples complicates fine-tuning approaches. In this study, we investigate the impact of positive and negative samples and their diversity and propose a controllable data synthesis pipeline, thus overcoming challenges.
Learning from Synthetic Data. The use of synthetic data [jordon2022synthetic, long2024llms, guo2024generative] has been studied across various computer vision tasks, such as classification [shipard2023diversity], segmentation [yang2023freemask], and object detection [feng2024instagen]. With the rapid consumption of existing data resources driven by the development of Large Language Models (LLMs) and VLMs, researchers aim for models to achieve further improvements using synthetic data [peng2024synthesize, li2024synthetic, liu2024synthvlm, huang2024modifyingdataaddressgraph]. Synthetic corpora are generated for instruction fine-tuning of large models. At the same time, large models leverage recaptioned datasets to enhance quality and improve CLIP [fan2023improving]. To better support the learning of VLMs, direct generation of text-image pairs is pursued [liang2024synth]. However, existing generation methods only consider class information [wang2024attributed]. These methods rely on simplistic prompt templates, attributes, and class information, either manually defined or LLM-generated [dunlap2023diversify, wang2024template]. To address the issue of insufficient data in VLM personalization scenarios, we supplement the dataset by synthesizing concept-centric positive and negative sample data. We adopt a unified framework to control the generated samples.
@@ 3 Method
We present a comprehensive data synthesis pipeline to address the challenge of personalizing Vision-Language Models (VLM). This pipeline views the concept information as a tree structure and controls the generation of positive and negative samples via tree operations. It also encompasses a well-designed filtering strategy to select the generated samples. The overall design of our pipeline is illustrated in Fig. 2. Specifically, in part 1, we systematically study the impact of positive and negative samples on VLM personalization tasks. Building on this, we introduce the CaT framework and related tree operations in part 2. Furthermore, to ensure the quality of the synthetic samples, we propose a filtering strategy based on the PCS score in part 3.
Figure 4: Concept tree synthesis and editing operations for sample generation. (a) A concept tree $T_{C}$ is constructed using the CaT framework and reference images. Positive samples are synthesized with a fine-tuned diffusion model guided by the root node $R(C)$. Easy negatives are generated by changing the class of $R(C)$, while diverse hard negatives are created by applying three editing operations to the original tree, using an unfine-tuned model. (b) Multiple concept trees can also be merged to support multi-concept personalization data generation. (c) The three operations are visualized: adding a “mood” dimension alters emotions; removing the “object” simplifies the scene; modifying a dimension changes behavior.
@@ 3.1 Impact of Positive and Negative Samples
In this section, we examine the effects of positive and negative samples respectively and their diversity on the personalization of VLM. Based on the positive samples provided by existing datasets and negative samples retrieved online, we conducted a series of experiments with the Yo’LLaVA method. As illustrated in Fig. 3 left, we found that an increase in positive samples consistently improves outcomes across all tasks. While easy negative samples primarily enhance recognition abilities, hard negative samples improve conversational capabilities. Interestingly, the captioning task fundamentally reflects conversational abilities. However, since the evaluation method assesses the presence of a concept identifier related to recognition capabilities, the training data that includes both easy and hard negative samples yields the best performance. Furthermore, we find that easy negative samples are essential. Using only hard negative samples can help the model perform well in conversational tasks, but recognition ability significantly declines. However, recognition tasks are the first step for VLMs acting as personal assistants, representing an essential capability that is crucial. Hence, easy negative samples are irreplaceable due to their diversity and broad sourcing, which offers valuable context for the model’s learning process.
Based on the above mentioned observation, we conducted additional experiments that examined different levels of diversity for negative samples. Here, diversity is calculated by clustering a set of negative samples and calculating the distance from the cluster center to each sample. From Fig. 3 middle and right, keeping the number of positive samples unchanged, it is evident that the gain curves for easy and hard negative samples differ across various tasks, indicating that their diversity requirements are distinct. Low diversity of both easy and hard negative samples can simultaneously restrict the model performance trained on them. When the diversity of hard negative samples is excessively high, it often introduces noise during training, resulting in a decline in model performance. These experiments clarify the role of diversity and motivate us to manage diversity when creating negative samples. Therefore, we propose representing the concept as tree to facilitate the generation of negative samples with various difficulties and diversities.
@@ 3.2 Concept as Tree Framework
Concept Tree Representation and Construction. To enable controllable data generation, inspired by SSDLLM [luo2024llm], we represent each concept $C$ as a hierarchical tree $T_{C}=(R(C),D_{C},A(D_{C}))$. As shown in Fig. 4(a), $R(C)$ serves as the root node, encapsulating the high-level category of the concept (e.g., “cat”, “dog”, etc.), while $D_{C}=\{D_{1},D_{2},\dots,D_{m}\}$ defines a set of attribute dimensions that distinguish various aspects (e.g., “appearance”, “behavior”, “location”, etc.). Each dimension $D_{i}$ is associated with a set of attributes $A(D_{i})$, collectively forming $A(D_{C})=\{A(D_{1}),A(D_{2}),\dots,A(D_{m})\}$ (for instance, the “behavior” dimension might include attributes like “sitting”, “lying” and “climbing”, etc.).
@@
Multi-Concept Scenarios. Our CaT framework seamlessly supports multi-concept scenarios by merging different concept trees according to the specific scenario, as illustrated in the Fig. 4(b). We generate training data for multiple concepts using the same approach as that used for single concepts. The scalability of the tree structure allows our framework to adapt flexibly to the continuous addition of new concepts, which is a significant advantage of our CaT framework.
Tree Editing Operations. We define three fundamental tree editing operations $F(R(C),D_{C}^{\prime})$ (see Fig. 4(c)), which modify $T_{C}$ to systematically generate negative samples:
@@
Figure 5: PCS-based filtering and visualization of high-quality image selection. (a) We apply patch shuffling to synthetic images and extract features $F_{o}$, $F_{d}$, and $F_{r}$ from original, disturbed and reference images. Image similarities $S_{o}$ and $S_{d}$ are computed to select images with high PCS scores. (b) Images rich in CS information are in green and those rich in CA information are in red. Cosine similarity can not distinguish between two types, but PCS score effectively filters out unqualified ones.
@@ 3.3 Sample Filtering with PCS Score
To ensure the quality of synthesized samples, we implement a filtering process for all generated samples. We classify the information contained in images into two types: concept-specific (CS) and concept-agnostic (CA). The former refers to features that are distinctive to the concept, such as appearance, while the latter encompasses other elements present in the image, such as background. We consider synthetic data predominantly containing CS information to be regarded as high-quality samples. Conversely, samples with more CA information may be categorized as low-quality samples.
Building upon this, we first mix patches from the reference and original images, subsequently calculating the similarity between the altered images and the reference images using the CLIP model (Radford et al., 2021). The difference in similarities before and after perturbation is defined as the Perturbation-based Concept-Specific (PCS) score. Typically, the PCS score of an image is expected to be high due to the alteration of patches, thereby disrupting the original Concept-Specific (CS) information. This disruption causes CLIP to struggle to discern the conceptual relationship between the synthetic image and the reference image, resulting in a significant decrease in similarity between the altered and reference images. However, if the similarity is based on Conceptual Attribute (CA) information rather than CS information, the similarity remains relatively unchanged, leading to a low PCS score. Therefore, we claim that the PCS score can be used to determine the in- formation components in the image. A high PCS score signifies a predominance of CS information, while a low PCS score indicates a predominance of CA information. We also conducted ablation experiments with alternative image encoders, presented in the Appendix, to demonstrate the generalizability of our method.
As illustrated in Fig. 5(a), let the reference image be $I_{\text{reference}}\in\mathcal{D}_{\text{user}}$, the synthetic image before perturbation be $I_{\text{original}}$ and the image after perturbation be $I_{\text{disturbed}}$. We extract the visual features $F_{r},F_{o}$ and $F_{d}$ of the corresponding images using CLIP [radford2021learning] image encoder $f_{\theta}$:
@@
The difference between $S_{o}$ and $S_{d}$ is the PCS score. Based on this score, we can select high-quality synthetic data, i.e., $G(I)=\{I\mid\text{PCS}(I)>\tau_{PCS}\}$, where $G(I)$ represents high-quality samples and $\tau_{PCS}$ is the threshold for containing more CS information. As shown in Fig. 5(b), when only using the filtering method based on cosine similarity, both the images in the green and red boxes achieve high similarity scores. However, the concept in the red box lacks specific visual features of the reference image’s concept, such as hair, facial structure and body shape. Instead, the background and other CA information in these images have a greater impact on the similarity. However, through the proposed perturbation-based filtering method, we can effectively identify and filter out images with more CA information using the PCS score, while retaining high-quality images with more CS information for better personalization.
@@
GPT-4o+Prompt 0.746 0.888 0.712 0.728 0.836 0.856 0.932 0.897 0.891 0.969
MyVLM(Real) 0.795 0.779 - 0.640 0.714 0.911 0.897 - 0.938 0.921
@@
RAP-MLLM 0.747 0.832 0.709 0.424 0.711 0.845 0.917 0.874 0.870 0.937
Table 1: Results of synthetic data used in different personalization methods. All methods achieve improvements across three datasets when using synthetic data. When trained with a combination of original positive data and all synthetic data, MC-LLaVA can even achieve results comparable to GPT-4o. Noting that the Choice-T is a text task, but MyVLM requires concept images to load the corresponding embeddings, making it unable to perform the task. The numbers following “+” represent the improvement of the new SOTA compared to original baseline. Real = Original Data; Syn = Synthetic Data; P = Plus.
4 Experiments
4.1 Experimental Setup
Implement Details. Regarding the experiments conducted on all datasets, the quantity of positive and negative samples for all the baselines follows their original settings. The training data used for Baseline (syn) is entirely generated by CaT. Baseline (Real+Syn) is articulated as comprising 1 to 3 original concept images and several synthesized positive samples, ensuring the same amount of the original positive samples. Baseline (Real+Syn) (Plus) indicates that an equal number of synthesized positive samples have been added to the previous positive samples. All negative sample quantities in the above experiments are strictly controlled for consistency. With regard to the synthesized images, we utilize GPT-4o to generate the instruction text pairs required for their training. Details of baseline, datasets and training hyperparameters can be found in the Appendix.